### v4

# VerifiableFL: Providing Verifiable Claims to Federated Learning using Exclaves

Abstract.
In federated learning (FL), data providers jointly train a machine learning model without sharing their training data. This makes it challenging to provide verifiable claims about the trained FL model, e.g., related to the employed training data, any data sanitization, or the correct training algorithm—a malicious data provider can simply deviate from the correct training protocol without detection. While prior FL training systems have explored the use of trusted execution environments (TEEs) to protect the training computation, such approaches rely on the confidentiality and integrity of TEEs. The confidentiality guarantees of TEEs, however, have been shown to be vulnerable to a wide range of attacks, such as side-channel attacks.
We describe VerifiableFL, a system for training FL models that establishes verifiable claims about trained FL models with the help of fine-grained runtime attestation proofs. Since these runtime attestation proofs only require integrity protection, VerifiableFL generates them using the new abstraction of exclaves. Exclaves are integrity-only execution environments, which do not contain software-managed secrets and thus are immune to data leakage attacks. VerifiableFL uses exclaves to attest individual data transformations during FL training without relying on confidentiality guarantees. The runtime attestation proofs then form an attested dataflow graph of the entire FL model training computation. The graph is checked by an auditor to ensure that the trained FL model satisfies its claims, such as the use of data sanitization by data providers or correct aggregation by the model provider. VerifiableFL extends NVIDIA’s NVFlare FL framework to use exclaves. Since exclaves require minor yet proprietary hardware changes, we implement a variant of our design with similar performance characteristics but slightly weaker security guarantees on existing AMD SEV-SNP hardware. We show that VerifiableFL introduces less than 12% overhead compared to unprotected FL model training.
1. Introduction
As machine learning (ML) models continue to be used in new and diverse applications, there is increasing impetus towards providing information about how these models were trained. For example, the EU AI Act (EUAIAct), which came into force in August 2025, requires providers of general-purpose ML models to provide “information on the data used for training, […] including the type and provenance of data and curation methodologies (e.g. cleaning, filtering etc.), the number of data points, their scope and main characteristics; how the data was obtained and selected as well as all other measures to detect the unsuitability of data sources and methods to detect identifiable biases, where applicable”.
In response to such requirements, datasheets for datasets (datasheetsdatasets) describe the composition of training datasets, and model cards (modelcard) describe model training. Providers of open-source models, such as the recently-trained Apertus (apertus), publish the code and datasets used for training; providers of closed-source models also make claims about their training process and datasets: e.g., the model card for OpenAI’s GPT-5 describes how the training data was pre-processed to remove personal information (gpt5). Such claims may be relevant to model providers as well as downstream users. For example, model providers may want to demonstrate to regulators that a model has not been trained on personal data or copyrighted content; alternatively, they may want to show that the model has been trained on certain data, such as under-represented languages, as claimed by Apertus (apertus).
At present, such claims about model training and datasets are made without proof and thus cannot be verified. For closed-source models, it is impossible to verify claims when interacting with the model via an inference API. Even for open-source models, it is generally infeasible to verify that a given model is indeed the result of executing the provided training code on a specified dataset—doing so would essentially require repeating the training process, which would incur substantial costs, and the trained model may still differ due to non-determinism. Today, model claims are therefore accepted based on the reputational trust in model providers.
In federated learning (FL) (mcmahan2017communication; kairouz2021advances; jere2020taxonomy) scenario, the situation becomes worse, because training is performed by (potentially many) data providers, rather than a single model provider. Claims about models trained using FL thus require reputational trust in all data providers, as well as the model provider. Nevertheless, FL is an appealing approach in many settings, because it allows data providers to retain control over datasets. Its use has been proposed for applications ranging from training large predictive language models on user input (hard2018federated; ramaswamy2019federated), to training fraud detection models while protecting transaction data of competing financial institutions (yang2019ffd).
In this paper, we explore the question whether it is possible “to provide verifiable claims about how a model was trained using FL by multiple data providers”. We observe that this would require systems support during training to discover any deviations by parties from the agreed training protocol, which would invalidate specific claims. For example, the system would have to discover if a data provider skips a given pre-processing step that removes personal information from a dataset. There could be various reasons for such deviations, ranging from accidental (e.g., bugs in the code) through to actively malicious (e.g., deliberately skipping steps to reduce costs).
A strawman design for creating an FL training system with verifiable claims about the trained FL model would be to execute the FL training computation inside trusted execution environments (TEEs), such as Intel SGX enclaves (arnautov2016scone) or AMD SEV-SNP confidential virtual machines (VMs) (sev2020strengthening). TEEs ensure that the integrity and confidentiality of computation is preserved by the hardware: they isolate the TEE memory used for the computation and offer attestation proofs, which contain a signed measurement of the TEE memory taken at deployment time. Indeed, prior designs of FL training systems have used TEEs to offer claims about trained ML models by leveraging such attestation proofs (openfl; quoc21secfl; guo2024trustworthy).
We argue, however, that TEEs are the wrong abstraction to create verifiable claims about FL model training for two reasons:
(1) Existing solutions focus on static attestation proofs. In FL training systems with TEEs (openfl; quoc21secfl; guo2024trustworthy), TEEs produce memory measurements at deployment time to provide a static attestation proof that establishes that the TEE runs the expected FL system and training code. A model provider can then check each data provider’s static attestation e.g., when establishing an initial TLS network connection (openfl). Such static attestation proofs, however, are hard to relate to claims about the final trained model: verifiable claims depend on how the training process was carried out at runtime but a static attestation proof only capture the state at the start of the training process. For example, if a data provider skips data sanitization in a given training round to poison the model, this would be difficult to discern from a single static attestation proof at deployment time.
(2) Existing attestation relies on confidentiality. As we show, TEE implementations rely on confidentiality to provide unforgeable attestation proofs. TEEs attempt to guarantee confidentiality by isolating TEE memory from external access. Unfortunately, this is challenging due to advanced attacks against TEE confidentiality, e.g., by exploiting side channels (foreshadow; zombieload; li2022systematic). If an attacker compromises the confidentiality of a TEE, they will have also compromised the integrity of its attestation proofs. For example, a distributed FL training system that relies on attested TLS connections between data providers and the model provider (guo2024trustworthy) would be vulnerable to an adversarial data provider who manages to leak the TLS session key from their TEE using a side-channel. This would allow them to deviate from the correct training protocol without being detected because they could forge the static attestation proof.
We observe that FL training should instead rely on a set of runtime attestation proofs about data transformations performed during FL training. The fact that particular data transformations, such as data sanitization (granite), local training (kairouz2021advances), poisoning mitigations (yin2018byzantine; flshield; gabrielli2023protecting) and model aggregation (yin2018byzantine), have been carried out faithfully can then be related to claims about the final trained FL model. Importantly, these runtime attestation proofs only require guarantees about the integrity of data transformations and not their confidentiality. Therefore, our approach can avoid the more challenging security model of TEEs, which must protect against both confidentiality and integrity attacks.
To realise the above approach, our work makes the following two contributions:
(1) We describe a new abstraction called an exclave, which is a hardware-based integrity-only execution environment. The security guarantees of exclaves do not depend on keeping their data secret from the exclave operator. Unlike TEE implementations, the integrity of exclaves is still preserved even if the exclave operator reads all data within the exclave, either directly or through a side channel. Exclaves support attestation through a hardware-based mechanism (e.g., a secure co-processor (sev2020strengthening)) to measure code integrity and sign an attestation statement at runtime that can include other runtime data. The signing keys for attestation are used and maintained by the hardware only, and are never accessible to software, making exclaves resilient to data leakage attacks. We show how exclaves can be realized through minor modifications to existing hardware-based TEE implementations.
(2) Using exclaves, we describe the design and implementation of VerifiableFL, a system for executing FL training jobs that provides verifiable claims about trained FL models. VerifiableFL uses exclaves to generate runtime attestation proofs of all executed data transformations in the FL job, covering the transformations performed by the data providers and the aggregations performed by the model provider in each training round. Each attestation proof, which we call an exclave data record (EDR), binds together through hashes: (i) the data inputs to the transformation; (ii) the code performing the transformation; and (iii) the resulting data output of the transformation. The EDRs are collected and used to construct an end-to-end exclave dataflow graph (EDG) to establish global properties. The EDG provides a complete and integrity-protected representation of all data transformations that led to the final trained FL model. It can be used by a third-party auditor to verify claims about the FL model, e.g., that data pre-processing and model training has been carried out faithfully by data providers across rounds and that the model provider has aggregated all model contributions by the data providers without bias.
We create a prototype implementation of VerifiableFL that extends the existing FL framework, NVFlare (roth2022nvidia), to execute individual FL tasks, which carry out data transformations in an FL training job, inside exclaves. VerifiableFL achieves this without requiring changes to the FL jobs or the framework implementation. Since our exclave design requires minor changes to proprietary TEE firmware, we implement a variant of our exclave design on AMD SEV-SNP using its virtual TPM feature, thus offering similar performance characteristics but slightly weaker security guarantees.
Our experiments show that VerifiableFL can provide verifiable claims such that the FL model was trained with the correct algorithms and datasets with low overhead: the execution of the FL training job within exclaves and the generation of runtime attestation proofs reduces training throughput by less than 12% compared to a baseline without exclaves.
2. Verifiable Claims in Federated Learning
Next, we provide background on training machine learning models using federated learning (§2.1) and discuss the challenges when providing verifiable claims about trained models (§2.2). After that, we introduce trusted execution environments (§2.3) and survey whether such approaches (or other techniques) can provide verifiable claims about trained models (§2.4). Finally, we consolidate all of the above into a precise threat model and set of security requirements for our desired solution (§2.5).
2.1. Federated learning
Federated learning (FL) enables the training of machine learning (ML) models while allowing data providers to retain control over their data. While FL applies to many scenarios, we consider the specific case of collaboration between multiple participants who wish to train models jointly using proprietary datasets, either for commercial gains, or to obtain models for their own use.
Figure 1. Overview of federated learning
As shown in Fig. 1, FL participants jointly train a model: data providers perform training locally using their own datasets and periodically send updates (either gradients or models) to a centralized model provider. The model provider aggregates the updates and distributes updated model versions to the data providers. This process repeats for a fixed number of rounds until the model converges.
Techniques exist to ensure the protection of datasets or the quality of the trained FL model. For example, data providers may apply differential privacy (DP) (dwork2014algorithmic) to bound the contribution of individual data records in the training dataset. In DP, a data provider adds noise to the model updates to protect the privacy of its dataset. In addition, to ensure the quality of the trained FL model, FL jobs may include algorithmic approaches to detect or mitigate against poisoning attacks (tolpegin2020data), e.g., using outlier detection (tolpegin2020data) or robust aggregation (blanchard2017machine; yin2018byzantine).
2.2. Claims about trained FL models
There is an increasing requirement to provide information about aspects of the training process for ML models. Proposals for providing such information, which are now being used in practice, include datasheets for datasets (datasheetsdatasets) and model cards (modelcard). Analogous to their use in electronics, the idea of datasheets for datasets is to describe the composition and intended use of a training dataset. Model cards perform a similar function, but for a trained model, as shown in the example in Fig. 2. Notably, model cards should ideally include information such as which training algorithm was used, which training datasets were used (e.g., by referring to the datasets’ datasheets), as well as details of any pre-processing performed on the datasets before training (e.g., to filter out certain types of content).
1 "model-information": {
2 "name": "...",
3 "model-provider": "...",
4 "architecture": "...",
5 },
6 "training-protocol": {
7 "dataset": "...",
8 "algorithm": "...",
9 "safety-mechanism": "...",
10 },
11 "benchmark-result": "...",
12 "intended-use": "...",
Figure 2. Example excerpt from a model card
In fully open-source models, the exact training code and dataset are usually made publicly available. In open-weight or closed-source models, the training code and dataset may be confidential, but model providers may still make public claims about the training code, dataset, and data pre-processing. For example, the card for OpenAI’s GPT-5 model claims that their data processing pipeline “includes rigorous filtering to maintain data quality and mitigate potential risks” and that they use “advanced data filtering processes to reduce personal information from training data” (gpt5). For all types of models, information about training may need to be provided to regulators. For example, Article 53 of the EU AI Act requires model providers to provide this information, “upon request, to the AI Office and the national competent authorities” (EUAIAct).
Making claims in FL training, e.g., that particular datasets or attack mitigations have been used, is even more challenging. Claims about a trained FL model require trust in the model provider and all data providers who contributed to the training. There could be various reasons for any of these participants to deviate from the agreed training protocol, either unintentionally or deliberately. For example, a data provider may accidentally omit or misconfigure steps in the pre-processing of the training data, leading to undesirable data being used for training. Alternatively, a data provider may promise to train on a particular dataset (e.g., patient data) but may only train on a subset of this data to reduce costs, resulting in a discrepancy between the claimed and actual training dataset.
It is an open question how to verify such claims about model training, even with access to the training code and datasets. For closed-source models that are only accessible via inference APIs, there is no way to verify claims about the model training; for open-source models, one could attempt to repeat the training process, but depending on the size of the model and the length of training, this may require significant hardware resources and thus incur substantial costs. Even if this were feasible, there is no guarantee that the trained model would match the original model, due to the inherent non-determinism in the training process. Therefore, claims about model training typically have to be accepted based on reputational trust in the entity making the claim. In an FL scenario, however, this reputational trust would have to extend to all participants, i.e., the model provider and also all data providers.
2.3. Trusted execution environments
Trusted execution environments (TEEs) try to ensure both the confidentiality and integrity of execution in a verifiable manner. They isolate the code and its data, preventing even privileged attackers, such as malicious administrators, operating system (OS) kernels, or hypervisors, from accessing the data or tampering with execution. TEEs also support attestation to enable external parties to verify a measurement taken over the TEE memory, which includes both the code and data. TEEs, including attestation, are implemented using hardware-based mechanisms, such as memory encryption, memory access checks, and hardware signatures for attestation.
TEEs exist at different granularities: ARM TrustZone (trustzone) provides a separate trusted world; Intel Software Guard Extensions (SGX) provides isolation at a process granularity using enclaves. More recently, VM-based TEEs, such as AMD SEV-SNP (sev2020strengthening), Intel TDX (mishra24tdx; cheng2023intel), and ARM CCA (cca24arm; li2023enabling), isolate entire VMs from the hypervisor and other privileged system software.
Hardware support for TEEs has also been integrated into GPUs. For example, NVIDIA’s H100 GPUs support a unified TEE, spanning the CPU-based TEE and the GPU (dhanuskodi2023creating; roth2022nvidia; gpublog). It protects sensitive GPU state, including GPU memory and configuration registers, from external access by the hardware, and all communication between the CPU-based TEE and the GPU is encrypted and integrity-protected. Graphcore IPUs provide similar TEE functionality (vaswani2023confidential).
2.4. Approaches for verifiable claims
Since TEEs can protect computation at a hardware level and provide attestation proofs, an FL model provider could require all data providers to perform their model training inside TEEs and use remote attestation to demonstrate this in a provable manner.
Existing approaches therefore propose the hosting of FL training jobs in TEEs: CFL (guo2024trustworthy) demonstrates how to host FL frameworks in TEEs based on SEV-SNP VMs coupled with NVIDIA Confidential GPUs; SecureFL (Kuznetsov2021SecureFL) is an FL training system that integrates SGX-based TEEs with ARM TrustZone on the edge to ensure privacy of both models and data; PPFL (mo_ppfl_2021) uses TEEs to prevent data providers from observing the model during training, thus preventing data reconstruction attacks; SecFL (quoc21secfl) includes a policy manager to enforce attestation policies agreed between all clients; OpenFL (openfl) uses SGX-based TEEs during FL execution to prevent model leakage; GradSec (valerio21gradsec) targets FL privacy threats by protecting a subset of the model layers inside TEEs; and CrowdGuard (rieger_close_2022) mitigates model poisoning by employing TEEs to perform client-side cross validation on local models.
Using TEEs to provide verifiable claims about a trained FL model, however, gives rise to two challenges: (1) the attestation proofs offered by TEEs are based on measurements of the TEE memory at deployment time. Such static attestation proofs includes a hash over the FL job code and the FL framework code. Typically they do not include a hash over the datasets, given that the data may be unavailable at deployment time. This makes it challenging to relate such an attestation proof to any claims about the final trained FL model: for an auditor to verify claims linked to such attestation proofs, they would have to re-run the FL training job and observe if they obtain the same final trained FL model from the same TEE attestation hashes.
In addition, (2) TEE attestation relies on the confidentiality of TEE memory, because existing attestation implementations use attestation keys maintained in TEE memory. Since TEEs guarantee the confidentiality of TEE memory, this approach is sound, unless the TEE confidentiality can be compromised. Here, TEEs have witnessed a long sequence of data leakage attacks, often by exploiting difficult to mitigate architectural side-channels (li2022systematic; wang2023pwrleak; spectre; li2021cipherleaks), which can be used to leak secrets from TEE memory. If such an attack is used successfully to leak the attestation key from a TEE, the integrity of attestation proofs can be compromised by an adversary. This would enable the adversary to produce bogus attestations, which they could use as evidence for false model claims.
Therefore, prior work has observed that guarantees about confidentiality and integrity should be separated in trusted execution. Sealed-glass proofs (SGP) (sgp) use SGX-based TEEs as “transparent enclaves” to provide attestation proofs of computation over secret inputs, while allowing a host to inspect the TEE memory. Transparent enclaves, however, still assume that attestation keys remain secret: with SGX enclaves, the attestation key is held inside a separate quoting enclave (QE) that resides in TEE memory, which makes it vulnerable to side-channel attacks (foreshadow).
Instead of relying on hardware mechanisms, other approaches have used algorithmic and cryptographic techniques to establish claims about trained models. Proof of learning approaches (jia2021pol; abbaszadeh2024zero) build on techniques such as proof-of-work (PoW) (pow), verified computation (verified_computation), and zero-knowledge proofs (zkp1; zkp2). These approaches, however, either incur prohibitively high overheads or can be subverted (zhang2022adversarialPOL; fang2023broken). For example, zero-knowledge proofs such as zkSNARK (zkp1) introduce orders-of-magnitude computational overheads (zkp2), making them impractical for FL training. PoW-based mechanisms (pow) require the majority of the participants to follow the protocol correctly. This assumption does not hold in FL training, because participants are mutually untrusted and may deviate from the correct training protocol.
2.5. Threat model and security goals
In our threat model, we assume that any participant in the FL training job (i.e., any data provider or the centralized model provider), or any set of colluding participants, could be adversarial. This also captures the case where one or more honest participants have been compromised by an adversary and thus act adversarially.
Adversary goal. The adversary aims to deviate from the predefined FL training protocol, in a way that modifies or influences the final FL model and for this deviation to remain undetectable by other participants and/or third-party auditors.
This broad adversary goal encompasses many specific attacks that different participants could attempt: e.g., an adversarial data provider may perform fewer rounds of training to reduce resource costs; or skip a pre-processing step that would otherwise filter out specific data; or change their dataset during training to bias the final model. Similarly, an adversarial model owner may omit updates from specific data providers during the aggregation step to bias the final model; or send different models to different data providers to infer information about a specific data provider’s private dataset (boenisch2023curious). More generally, any adversarial participant may want to deviate from the predefined FL training protocol by changing their training code or by modifying or falsifying any data communicated during the FL training job.
Adversary capabilities. In an FL scenario, we assume a strong adversary with privileged access on a participant’s system. This includes full control of the OS and hypervisor of a participant, as well as the ability to intercept and modify all inbound and outbound network communication. The adversary can carry out side-channel attacks to leak the contents of a participant’s memory, even memory that is protected by a TEE. We assume, however, that the adversary will not carry out advanced physical attacks against the hardware, such as physically extracting keys from an on-chip security co-processor, due to the high costs and expertise required.
Security goals. VerifiableFL aims to ensure that the aforementioned adversary is unable to deviate from the predefined FL training protocol, in a way that modifies or influences the final model, without detection. VerifiableFL thus aims to enable any participant or third-party auditor to verify claims about the FL training process, as described in §1.
Specifically, we define a verifiable claim to be any statement about the FL training process that could have been made by a (hypothetical) trusted omniscient observer with full visibility of the training process. Claims could include statements about the executed code for training and aggregation, the number of participants and rounds of training, or any other runtime property of the FL job. It is also possible to make statements about the precise data used for training, but sometimes it is desirable for the data to remain private to the data providers. In such cases, statements can be made about a representation of the data, e.g., a blinded hash to confirm that the data has not been changed during the training process. In addition, claims may require statements about what preprocessing (e.g., filtering) has been run on the data, based on the code used.
Security non-goals. Given our assumed adversary capabilities, we cannot prevent deviation from the protocol, but we require that any such deviation is detectable. Similarly, we cannot guarantee the availability of verifiable claims, as there are various ways in which our strong adversary can prevent such claims from being made (e.g., through denial-of-service attacks). We require that, if a claim is made and verified, it must be a faithful representation of the FL training job. Finally, we do not aim to detect poisoning of a training dataset before training begins. There is a large body of related work on the detection or prevention of such attacks (rieger_close_2022). Instead, VerifiableFL can support such mitigations by (i) ensuring that they have been applied correctly by the respective participants; and (ii) allowing third parties to verify that these mitigations have been in place during training.
3. Exclaves
In this section, we introduce exclaves, a new abstraction that we use to overcome the challenges described above. We first describe the exclave concept (§3.1) and then discuss how exclaves can be implemented (§3.2).
3.1. Exclave concept
Current TEE implementations use attestation to establish a trust relationship between the TEE and a verifier. For example, an attested TLS network connection (fossati-tls-attestation-08) can prove that the verifier communicates with the TEE and that all messages received via that TLS connection are sent by the TEE.
However, as explained in §2.5, an adversary who has the ability to leak data from a TEE (e.g., through a side channel) can undermine the security guarantees of current TEE implementations, including those intended to secure FL training. In the above example, this adversary could leak the session key of the attested TLS connection and then masquerade as the TEE. In the presence of such an adversary, we cannot rely on an approach that requires data within the TEE to be kept secret from the adversary.
Exclaves. Therefore, we introduce an exclave, which is an integrity-only execution environment that does not contain data that must be kept secret from the adversary. Similar to TEEs, an exclave ensures that code is executed without external interference, and that the data within the exclave can only be modified by code within the exclave. Thus, the exclave offers integrity guarantees for its code and data, even against privileged software running on the same host.
In contrast to TEEs, however, exclaves do not offer confidentiality guarantees—the only piece of data that must be kept secret from the adversary is a single attestation signing key. As we explain in §3.2, this key can be held by a secure coprocessor, which is hardened against side-channel attacks by being physically separate (frazelle2019securing).
Code running within the exclave can then request the coprocessor to produce a signed attestation report, which contains (i) a measurement of the exclave’s code and data at deployment time and (ii) additional user-defined data supplied by the code inside exclave at runtime. This type of attestation report can be used to prove that a particular statement was made by the exclave.
Exclave Data Records (EDRs). The attestation report can be used to produce an unforgeable record of data transformations carried out by the exclave (i.e., any computation inside the exclave that takes a set of data inputs and produces a set of data outputs). Exclaves can thus be used to attest individual data transformations at runtime. By performing a data transformation within an exclave, the exclave can produce a runtime attestation proof that binds the data transformation to the specific code and data of the transformation. We refer to this attestation proof as an exclave data record (EDR).
Formally, an EDR is an existence proof that at least one execution of the specified code produced the specified data transformation. In terms of its implementation, an EDR is signed statement that contains (i) a measurement of the exclave’s code, which defines the computation; and (ii) a representation of the data transformation, which contains two key/value structures: one for inputs and the other for outputs. The keys are unambiguous names for each of the inputs and outputs (e.g., model_weights or training_data). The values are fixed-size cryptographic hashes of the actual inputs and outputs. The measurement of the exclave’s code is a cryptographic hash of the code pages that were loaded into the exclave at deployment time. Given access to that code, a verifier can thus recompute this hash and compare it against the measurement in the EDR. This is the same type of code measurement performed by TEEs.
3.2. Exclave implementation
An exclave can be implemented with minor modifications of a TEE implementation, such as AMD SEV-SNP (sev2020strengthening), by reusing its existing isolation and hardware-based attestation. For example, exclaves can rely on the memory isolation enforced by the memory management unit (MMU) in AMD SEV-SNP: on each address translation, the MMU checks the access policy stored in the reverse map table (RMP) and maintained by the secure processor, i.e., the platform security processor (PSP) (buhren2019insecure). In addition, the exclave uses the PSP to generate hardware-signed attestation reports at runtime. The key used for signing attestation reports, the versioned chip endorsement key (VCEK) (buhren2019insecure), is stored and managed by the PSP only.
However, exclaves must generate runtime attestation reports by passing data securely to the PSP, without relying on shared software secrets, as this would violate their secret-free nature. Currently, AMD SEV-SNP assumes that a TEE maintains a VM platform communication key (VMPCK) to protect communication with the PSP. An exclave implementation therefore requires a modification of AMD SEV-SNP to protect the communication with the PSP through an MMU access check instead of a shared secret: the exclave must rely on a special protected page, accessible only by the exclave and the PSP, to generate attestation report securely.
Implementing exclaves using AMD-SEV SNP thus would require the following firmware changes: since an exclave would be a guest CVM running on the host, the proposed protected page could be implemented as a guest-writable analogue of SNP’s existing context page, which is RMP-protected, PSP-writable, but only guest-readable. This would require: (1) defining a new “protected” page state whose RMP entry has the same metadata as the context page, but with ownership metadata set to the exclave’s ASID; and (2) extending the access control policy to allow writes from the exclave to the protected page when the ASID matches; and (3) adding a new SNP_PROTECTED_CREATE command to provision this protected page on exclave creation (similar to context page provisioning). With these changes, access to the protected page would be enforced by SNP’s RMP check, thus enabling secret-free attestation. These firmware changes to the SNP would be feasible, because they could reuse much of the existing implementation of the context page protection.
For other TEE technologies, achieving “secret-free” attestation would require the removal of privileged software components, such as a security monitor (SM), from the runtime attestation process. For example, the Keystone TEE (keystone) would need integration with a secure hardware signer (e.g., HSM/TPM (tpm)), and its SM would become an untrusted proxy for attestation requests to that hardware signer. Compared to other TEEs, AMD SEV-SNP already provides hardware-signed runtime proofs by the PSP, which matches an exclave’s requirements.
Since we are unable to modify the proprietary firmware in AMD SEV-SNP to implement the above change to key sharing, we emulate equivalent signing functionality for exclave attestation. As we describe in §6.3, our exclave implementation on unmodified AMD SEV-SNP hardware emulates the modified attestation approach using a virtual TPM as a stand-in for the PSP’s attestation functionality with comparable performance.
4. VerifiableFL Design
Figure 3. Design of VerifiableFL (Exclaves are shown in green.)
Next, we describe the design of the Verifiable Federated Learning (VerifiableFL), a system that produces verifiable claims about the training process and the resulting FL model. The goal of the VerifiableFL design is to use the abstraction of exclaves to produce runtime attestation proofs, in the form of EDRs, of all executed data transformations that led to the creation of a final trained FL model. The collected set of EDRs provides a trustworthy end-to-end record of the training progress across the rounds of the FL computation. Therefore, an auditor can use the EDRs to discover violations of specific claims about the trained FL model (see §5).
The use of exclaves in the design of an FL training system introduces two challenges: (i) the integrity of the produced claims must not rely on the confidentiality of the FL training process because, unlike TEEs, exclave execution is visible to individual FL participants; and (ii) to keep the TCB small, the produced claims should only require the minimum amount of code to be executed (and thus integrity-protected) inside exclaves. In particular, the integrity of code belonging to the FL framework should not have to be protected, as long as the EDRs that attest the data transformations also establish its correct execution.
As shown in Fig. 3, VerifiableFL executes an FL training job as a sequence of FL tasks that define the data transformations, which data providers and the model provider must execute in each training round. An FL framework such as NVFlare (roth2022nvidia) provides the task implementation, orchestrates task execution and provides APIs for data movement between tasks. VerifiableFL executes each task within an exclave, which ensures execution integrity and produces EDRs for each executed data transformation.
In a typical FL job, the tasks executed by data providers may include: (i) a local training task that takes a global FL model, reads a dataset from storage, computes a local model update and returns the updated local model; and (ii) a differential privacy (DP) task that takes a local model update, adds DP noise to protect its privacy, and outputs a DP-protected local model update. A model provider may execute: (iii) an aggregation task that receives DP-protected local model update from data providers and outputs an aggregated model; and (iv) a model update task that converts the aggregated model to a global FL model. The updated global FL model is sent to the data providers ahead of the next training round. Note that the rest of the FL framework executes outside of exclaves, thus keeping the TCB restricted to the data transformations that impact the trained FL model.
Since VerifiableFL executes tasks inside exclaves, the produced EDRs create a trustworthy record of the executed data transformations. For example, an EDR produced by the aggregation task contains hashes of (i) the local model updates received by the task in a round and (ii) the output aggregated model; an EDR produced by the model update task has hashes of (i) the aggregated model and (ii) the updated global model; and an EDR produced by the local training task contains hashes of (i) the received global model, (ii) the training dataset and (iii) updated local model. When the FL job is executed faithfully by all participants, the EDRs must show that the correct code for the task implementations has been executed. In addition, the hashes in the EDRs that denote the data inputs and outputs must match between EDRs executed by different tasks within a training round: e.g., the output hash of the updated global model in the EDR generated by the model update task executed by the model provider must match the input hashes of the global model in the EDRs generated by the local training tasks executed by the data providers.
Since the local training task reads a dataset from storage, VerifiableFL must also ensure that its input hash in the EDR remains valid. Therefore, a data provider must provide the training dataset as an integrity-protected disk image (dmverity) that uses a Merkle tree (merkletree) to ensure the integrity of the blocks in the disk image. When the dataset is accessed by the local training task in the exclave, VerifiableFL checks the dataset blocks against the Merkle tree: the root hash of the Merkle tree (combined with a random salt) serves as a non-revocable data commitment to the training dataset, which is included in the EDR.
When exclaves produce EDRs, the EDRs are sent to an EDR storage service. The EDR storage service maintains all generated EDRs during the execution of the FL job and makes them available for auditing. Note that the EDR storage service itself does not need to be trusted: since EDRs are integrity-protected by a cryptographic signature, any tampering would be detected. Similarly, if individual EDRs are withheld by an adversary from the EDR storage service, the introduced inconsistencies between EDRs (see below) would reveal this.
As we describe in §5, an auditor can now check the EDRs for any mismatches or inconsistencies, which would indicate a deviation from the faithful execution of the FL job, thus invalidating claims about the trained FL model. The auditor can be an FL participant, a consumer of the trained FL model who needs to verify correct training, or a regulator who audits the training process for compliance.
5. Auditing of Verifiable Claims
To verify that the data transformations led to the final trained FL model, VerifiableFL provides a mechanism to audit EDRs. For this, it constructs a graph based on all EDRs (§5.1). Using the graph, an auditor verifies claims about the FL job (§5.2).
5.1. Constructing an exclave dataflow graph
Figure 4. Example of an Exclave Dataflow Graph (Edges show reverse dataflow; nodes show Training, DP, Aggregation, and Model update tasks, respectively.)
The EDRs from all participants (collected by the EDR storage service) are used to construct a directed acyclic graph, which we refer to as the exclave dataflow graph (EDG). Fig. 4 shows a sample EDG: each vertex $v\in V$ is an EDR, which is defined by the tuple of $(\mathit{in},\mathit{code},\mathit{out})$ that refers to the data input, code and data output hashes from that exclave execution. Each edge $e\in E$ represents the (reversed) data flow between exclaves; it points from the downstream (consumer) vertex $v_{y}$ to the upstream (producer) vertex $v_{x}$:
(1) $e=\{(v_{y},v_{x})\mid v_{y},v_{x}\in V\text{ and }v_{y}.\mathit{in}\cap v_{x}.\mathit{out}\neq\emptyset\}$
The edge label represents the data that flows between the two exclave executions. It is defined as the non-empty intersection of the outputs of $v_{x}$ and the inputs of $v_{y}$ (since exclaves may have multiple inputs and outputs for a single execution):
(2) $l(v_{y},v_{x})=\{v_{y}.in\cap v_{x}.out\}$
We define the $\mathit{EDG}=(V,E,l)$ as a collection of vertices $V$, reversed dataflow edges $E$, and an edge label function $l$. For auditing, Alg. 1 shows how VerifiableFL constructs an EDG from a collection of EDRs $v$ that are attested as $\sigma(v)$. For each EDR, VerifiableFL first verifies the attestation with the hardware manufacturer, e.g., using AMD’s Key Distribution System (amd_kds) (line 1). If the verification succeeds, VerifiableFL adds the EDR to the EDG as a vertex; if it fails, e.g., due to an invalid signature or other tampering, the vertex is ignored (see Fig. 5(a)).
VerifiableFL then adds edges by iterating over all vertices. [^1] In practice, the structure of the FL job can be used to ascertain which vertices should be connected, avoiding naive iteration. If the inputs of an EDR vertex match the outputs of another, an edge is added between the two vertices, labeled with the dataflow between the exclaves (lines 1013). If the data was modified outside the exclave, e.g., by an adversary during transit, there is no match and no edge is added due to the mismatch between inputs and outputs (see Fig. 5(b)).
Algorithm 1 Constructing an EDG for auditing
0:  Collection of vertices $v\in V^{\prime}$, attestation reports $\sigma(v)\in\Sigma$
0:  $EDG=(V,E,l)$
1:  $V\leftarrow\{v\mid\mathit{verify\_attestation\_report}(\sigma(v))=\mathit{true}\}$
2:  $EDG\leftarrow\{V,E=\emptyset,l=\emptyset\}$
3:  // Step 1: Index vertices by input
4:  for each $\{v_{y}\in V\}$ do
5:    for each $\{input\in v_{y}.\mathit{in}\}$ do
6:      append $v_{y}$ to $vertexOf[input]$
7:  // Step 2: Create edge when input and output match
8:  for each $\{v_{x}\in V\}$ do
9:    for each $\{output\in v_{x}.\mathit{out}\}$
10:      if $\{vertexOf[output]\neq\emptyset\}$
11:       for each $\{v_{y}\in vertexOf[output]\}$
12:        $e\leftarrow(v_{y},v_{x})$
13:        $l(e)\leftarrow v_{y}.\mathit{in}\cap v_{x}.\mathit{out}$
14:        $E\leftarrow E\cup\{e\}$, $l\leftarrow l\cup l(e)$
15:  return $(V,E,l)$
5.2. Checking verifiable claims
(a) Invalid attestation
(b) Modifying data
in transit
(c) Incorrect orchestration
(d) Modified dataset
(e) Non-compliant dataset
with Sanitization (zoom in)
Figure 5. Variations of the EDG from Fig. 4 with deviations
The constructed EDG is audited to verify claims about the trained FL model. Next, we describe possible verifiable claims using examples of specific training-time deviations and explain how an auditor detects these deviations:
Claim 1: Correct FL training algorithm implementation. Any participant could modify the code of tasks before deployment. For example, a data provider could modify the training and DP tasks to skip data integrity checks, ignore certain training data, inject too much noise, or modify model updates. A model provider could modify the aggregation or model update exclaves to omit inputs, change the aggregation algorithm, modify model weights, or distribute different models to different participants. The modified code would be measured by exclaves and recorded in the EDR. Since the auditor traverses the EDG and checks each vertex against an allowed list of acceptable code measurements for tasks in the FL job, they detect the claim violation.
Claim 2: Correct transmission of training data. A malicious participant could modify the data in transit, either between exclaves (Fig. 5(b)), or when sent by the FL framework between participants. This would result in a mismatch between EDR outputs and inputs, thus missing an edge in the EDG. Since there would be a vertex with an input but no output edge, the auditor would identify the claim violation.
Claim 3: Correct use of DP/unbiased data aggregation. As shown in Fig. 5(c), a malicious participant could bypass the DP task execution and send the output of the training task directly to the model provider without DP noise. Similarly, a model provider could introduce bias by dropping contributions from certain data providers. This is more subtle than the preceding attacks, because it still results in a complete EDG. The shape of the resulting EDG, however, would differ from that of the expected EDG: it would miss vertices/edges, or have edges that span multiple training rounds, which would be detected by the auditor.
Claim 4: Integrity of training dataset. A malicious data provider could add, remove, or modify its dataset after training has commenced (e.g., performing an adaptive poisoning attack (a3fl)). If the data is modified while being read by an exclave, the verification against the Merkle tree fails; if the dataset is modified between training rounds (see Fig. 5(d)), the hash of the modified dataset is recorded in an EDG, enabling the auditor to discover the dataset change.
Claim 5: Appropriate dataset sanitization. A data provider may use a non-compliant dataset, e.g., with harmful content or personal information, without sanitization. If the FL job includes a sanitization task, the correct execution of data sanitization is recorded by the EDRs. Therefore, an auditor can compare the training tasks’s input dataset commitment against the set of data commitments produced by the sanitization task (Fig. 5(e)). Note that the auditor does not require access to the plaintext dataset to detect whether the dataset was exposed to a given sanitization task.
6. VerifiableFL Implementation
Figure 6. VerifiableFL implementation
In this section, we describe the VerifiableFL implementation (§6.1) and its components (§6.2), as well as our implementation of exclave attestation (§6.3) and integrity-protected storage for data (§6.4).
6.1. Overview
A goal of the VerifiableFL implementation is to be compatible with different FL frameworks and to support existing FL frameworks with minimal code changes. To achieve this, VerifiableFL adopts an orchestrator/worker model in which an untrusted FL framework still orchestrates the FL job and tasks execution inside exclaves.
Fig. 6 shows how VerifiableFL integrates with an FL Framework: the model provider submits an FL job description to the FL Framework  1 . The FL Framework orchestrates the job according to its description, and allocates FL tasks to the associated exclaves via Exclave Connectors  2 , which are VerifiableFL components responsible for creating and invoking exclaves. When an exclave receives a task, an Exclave Manager  3 forwards the task payload (e.g., global model) to a Task Executor. During task execution  4 , the task can access persistent storage via a Secure Storage module  5 , which provides integrity protection for on-disk datasets. When the task completes, a Runtime Attestor  6 attests the EDR of the exclave, and returns the execution result (e.g., local model) to the FL Framework.
Table 1. FL tasks deployed in exclaves
Exclave Task Description
Training StartTrain(...) Receive global model
GetWeight(...) Return local model
DP AddNoise(...) Perform local DP
Aggregation Accept(...) Receive Local Models
Aggregate(...) Aggregate local models
Model Update ModelUpdate(...) Weight diff to full model
6.2. Components
VerifiableFL has the following components:
FL Framework. VerifiableFL uses NVFlare (roth2022nvidia), which is an open-source, distributed FL framework by NVIDIA. It provides FL task implementations for many FL algorithms, e.g., federated averaging, federated optimization, and secure aggregation, and other algorithmic building blocks, e.g., model trainers, model aggregators, and task filters. Tab. 1 shows a set of tasks used in a typical FL job. VerifiableFL executes all of these FL tasks inside exclaves.
Exclave Connectors are gRPC clients (grpc) that NVFlare uses to execute FL tasks in exclaves. They are implemented as NVFlare tasks, which makes them compatible with the NVFlare API. To add an exclave, VerifiableFL includes the corresponding Exclave Connector in NVFlare’s job description file. This approach of using Exclave Connectors avoids modifications to the NVFlare code.
Exclaves. VerifiableFL implements exclaves as confidential VMs (CVMs) based on AMD SEV-SNP (sev2020strengthening). Within the CVM, exclaves are implemented as gRPC servers that execute FL tasks.
Exclave Managers expose exclave interfaces to NVFlare to support the orchestration of FL jobs. Exclave Managers receive FL tasks from NVFlare, deserialize their payloads, and forward the payload to the Task Executor. They then take the execution results from the Task Executor and pass them to NVFlare.
Task Executors execute tasks inside exclaves. A task Executor receives FL tasks from the Exclave Manager, loads the necessary data/code, and executes it. For example, the Task Executor for large language model (LLM) training initializes the task execution by loading the LLM model, the training code, and the training dataset. VerifiableFL reuses specific FL task execution code from NVFlare with minimal changes.
Runtime Attestor. The Runtime Attestor constructs EDRs and interacts with a secure processor for attestation. During task execution, it monitors the storage accesses and requests/responses passing through the Exclave Manager to capture exclaves’ input and output data. The Runtime Attestor records a hash of each data item together with metadata that describes both the intended usage (e.g., input or output) and properties (e.g., model or dataset). Once an input/output pair is collected, it constructs the EDR as a nested JSON object: keys contain the usage and properties of the data, and the values contain data hashes. The Runtime Attestor invokes the secure processor’s signing API with a SHA-256 digest of the EDR. The secure processor signs it together with a system measurement, and returns the signed EDR.
6.3. Exclave emulation
To implement exclave attestation without relying on software secrets, we emulate the attestation path using a virtual Trusted Platform Module (vTPM) as the secure processor in AMD SEV-SNP.
The vTPM provides attestation primitives, including measurement and signing, and executes at the highest privilege level in the CVM. Its signing key is endorsed by the PSP when the CVM launches. We use virtual machine privilege modules (VMPL) to create privilege levels: the vTPM runs in VMPL0 (highest privilege), while the exclave runs in VMPL2 (lower privilege). The MMU and RMP enforce isolation so that the exclave in a lower privilege level cannot interfere with the attestation process at higher privilege.
Attestation then uses an MMU-protected shared memory page, which is a VMPL2 page writable by both the vTPM and the exclave. For attestation, the vTPM reads the runtime claim’s hash from the shared page, extends it into the PCR23, signs the PCRs, and writes the signed report back.
6.4. Secure storage with commitment
The Secure Storage module provides integrity-preserving and verifiable data storage for exclaves, which is used to protect both datasets and task code.
Datasets. VerifiableFL uses Linux’s dm-verity (dmverity) for integrity protection, which checks accessed disk blocks against a Merkle tree. Data providers use the dm-verity tool to package datasets with a Merkle tree. The generated data commitment becomes a unique measurement over the dataset, which is included in EDRs by the Runtime Attestor. When setting up an exclave, VerifiableFL mounts the data disk and checks the expected data commitment against the Merkle tree.
Code. VerifiableFL also uses dm-verity protection for the exclave’s code image by using the Merkle tree root hash as a code measurement. When an exclave is launched, the vTPM includes the code measurement in the PCR11 field of the attestation report. Since it uniquely identifies the exclave code, it can be used by the auditor to determine whether an FL task was executed by the correct exclave.
7. Evaluation
We evaluate VerifiableFL to answer the following questions: (i) what is VerifiableFL’s end-to-end training performance compared to existing solutions? (§7.2); (ii) how does VerifiableFL scale with an increased number of data providers and GPUs? (§7.3); (iii) what is the performance breakdown of VerifiableFL? (§7.4); (iv) what overhead does VerifiableFL’s runtime attestation add? (§7.5); (v) what is the impact of secure storage on throughput and latency? (§7.6); and (vi) what is the auditing overhead as VerifiableFL scales? (§7.7)
(a) Language modeling
(b) Named-entity recognition
(c) Image classification
Figure 7. End-to-end training throughput
7.1. Experimental setup
Testbed. We deploy confidential VMs (CVMs) on Azure. CVMs without GPUs run on AMD EPYC 7763v CPUs; CVMs with GPUs run on AMD EPYC 9V84 CPUs, both with SEV-SNP and vTPM enabled. CVMs with GPUs have one NVIDIA H100 GPU with Hopper Confidential Computing extensions (h100cc) enabled.
On Azure, we deploy 5 CVMs (EC4as_v5) to execute 1 model update exclave and 4 NVFlare clients; 5 CVMs (EC8as_v5) to execute 4 DP exclaves and 1 NVFlare server; and 1 CVM (EC16as_v5) as the aggregation exclave. Training exclaves are deployed on 4 CVMs with CPUs (DC32as_v5) or 4 CVMs with GPUs (NCC40ads_H100_v5). Each training exclave mounts an Azure P30 data disk, which has up to 200 MB/s of throughput and support 5,000 IOPS.
VerifiableFL implementation. We implement VerifiableFL in C++ and Python. It uses PyTorch v2.5.0 with CUDA 12.4 for the FL tasks, Azure Guest Attestation Library v1.0.2 for CVM attestation, and NVFlare v2.2.1 (roth2022nvidia) as the FL framework. [^2] Our open-source artifacts will be released upon acceptance. As described in §6.3, we emulate the attestation behaviour of exclaves using a vTPM implementation (vtpm) running on the AMD SEV-SNP platform.
Baselines. We compare VerifiableFL with the two baselines: (i) NVFlare (No security) executes the FL workflow with the NVFlare server and clients in regular VMs. Each NVFlare node runs in a monolithic VM with SEV-SNP features disabled; and (ii) NVFlare (CVM) (guo2024trustworthy) hosts each NVFlare client, as well as the server, in its own monolithic CVM. For a fair performance comparison, both NVFlare (CVM) and VerifiableFL use the vTPM for attestation.
Datasets. We use 3 public datasets: (i) Wikitext103 (merity2016pointer) consists of raw text with over 100 M tokens from Wikipedia articles; (ii) NCBI-disease (dougan2014ncbi) is a disease text mining dataset with 793 fully annotated PubMed abstracts; and (iii) CIFAR10 (krizhevsky2009learning) has annotated $32\times 32$ pixel images with 50,000 training and 10,000 testing samples.
ML models. We use 6 widely-used ML models, covering three different categories: (i) decoder-only transformers—we use GPT-2 (radford2019language) with 124 M parameters, and a distilled version of GPT-2, DistilGPT2 (sanh2019distilbert), with 82 M parameters; (ii) encoder-only transformers—we employ BERT (kenton2019bert) with 110 M parameters, and MobileBERT (sun2020mobilebert) with 25 M parameters; and (iii) convolutional neural networks (CNN)—we use LeNet (lecun1998gradient), a 5-layer CNN, and VGG9 (simonyan2014very), a 9-layer CNN.
(a) GPT2
(b) BERT
(c) VGG9
Figure 8. Throughput with different numbers of clients
7.2. End-to-end throughput
We evaluate the end-to-end performance of VerifiableFL in terms of throughput. We run VerifiableFL with 1 FL server and 4 FL clients. Each client performs model training with either a CPU or GPU CVM, depending on the workload. After each training round, FL clients use the sparse vector differential privacy technique (SVT-DP) (lyu2016understanding) to add DP noise to the locally trained model. VerifiableFL runs each FL task for 10 rounds with federated averaging (FedAvg) aggregation. We measure the performance of VerifiableFL for our three ML categories: (i) language modeling; (ii) named-entity recognition; and (iii) image classification. Fig. 7 shows the measured end-to-end throughput results.
Language modeling. In this experiment, each client has a training exclave that executes the language modeling task in a GPU CVM (40 vCPUs; 320 GB RAM; 1 NVIDIA H100 GPU). We train both GPT-2 and DistilGPT2 models on the Wikitext dataset, which is partitioned into four subsets, one for each client.
Fig. 7(a) shows the training throughput. With DistilGPT2, VerifiableFL has a $6\%$ overhead compared to the NVFlare (No security) baseline; NVFlare (CVM) has a $1\%$ overhead. With GPT-2, when compared to NVFlare (No security), VerifiableFL and NVFlare (CVM) incur less than $10\%$ and $4\%$ overhead in training throughput, respectively. We observe that, in both cases, VerifiableFL is slower than NVFlare (CVM): in VerifiableFL, exclaves introduce runtime overhead due to the runtime attestations; and the distributed deployment of exclaves introduces network overhead when (de)serializing the ML model and transmitting the serialized ML model.
Named-entity recognition. Training exclaves execute in CPU CVMs (32 vCPUs; 128 GB memory). We use the MobileBERT and BERT models on the NCBI-disease dataset, which is split into four shards as each client’s local dataset.
Fig. 7(b) shows the training throughput. With MobileBERT, VerifiableFL and NVFlare (CVM) are $5\%$ and $4\%$ slower than NVFlare (No security), respectively. With BERT, VerifiableFL incurs a $9\%$ overhead compared to NVFlare (No security), and NVFlare (CVM) has a $4\%$ performance cost. We conclude that, when training MobileBERT, VerifiableFL has less performance impact compared to the BERT training. Since MobileBERT is $4.3\times$ smaller than BERT, it results in less network communication between exclaves in VerifiableFL.
Image classification. We also execute the image classification task with CPU CVMs and the same setup as before. Here, we use the LeNet and VGG9 models with the CIFAR10 dataset. Since the CIFAR10 dataset is relatively small, we replicate it four times and shard it into four subsets, ensuring each client receives sufficient training data.
Fig. 7(c) shows the performance results. For VGG9 training, VerifiableFL and NVFlare (CVM) incur $9\%$ and $4\%$ of overhead, respectively, when compared to NVFlare (No security). Compared to other workloads, VerifiableFL incurs a higher overhead for LeNet training: it is $11\%$ slower than NVFlare (No security). Even with NVFlare (CVM), LeNet training overhead is 7%. The significant CVM overhead with VerifiableFL’s extra scheduling and communication impact performance.
For moderate and large models (i.e., VGG9, MobileBERT, BERT, DistilGPT2, and GPT-2), VerifiableFL’s overhead is below $10\%$ compared to NVFlare (No security), which is slightly higher than NVFlare (CVM)’s overhead of $4\%$. Across all models, the overhead stays low. LeNet has a slightly higher overhead than other models, because its small model requires little training time, which makes it sensitive to the CVM overhead and VerifiableFL’s runtime and I/O costs.
7.3. Scalability
Next, we investigate VerifiableFL’s scalability. We increase the number of FL clients and measure VerifiableFL’s performance in terms of training throughput. Note that adding FL clients results in a corresponding increase in the degree of data parallelism. We use the same testbed settings and FL tasks as in §7.2.
First, we measure VerifiableFL’s scalability with the GPT-2 model for language modeling (Fig. 8(a)). Here, we scale VerifiableFL from 1 to 4 clients, each with a GPU. With $3\times$ increase in FL clients, VerifiableFL, NVFlare (CVM) and NVFlare (No security) have a $\times, 2.8×2.8\times$, and $2.9\times$ performance increase, respectively.
We also consider the BERT model for named-entity recognition (Fig. 8(b)), and the VGG9 model for image classification (Fig. 8(c)). In these experiments, we increase the number of CPUs when scaling the number of clients. With BERT, a $3\times$ increase in FL clients leads to a $2.8\times$, $3.0\times$, and $3.0\times$ increase in throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security), respectively. For VGG9, we observe that the throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security) increases by $2.9\times$, $3.0\times$, and $3.0\times$, respectively, when we quadruple the client count. Both results show that VerifiableFL scales effectively with more FL clients, similar to other baselines. Since all clients run in parallel, the runtime and I/O overheads of clients do not accumulate.
Table 2. Breakdown of execution time
Component GPT2 VGG9
time (sec) time (sec)
Training Exclave 8,659 (73%) 4,641 (97%)
DP Exclave 1,176 (10%) 21 (¡1%)
Aggregation Exclave 50 (¡1%) 1 (¡1%)
Model Update Exclave 29 (¡1%) 1 (¡1%)
FL Framework 1,983 (17%) 124 (3%)
7.4. Performance breakdown
We also provide a breakdown the execution time of VerifiableFL’s components when running the FL job from §7.2. Tab. 2 shows the results. With the GPT-2 and VGG9 training tasks, the training exclave dominates execution time (97% for VGG9; 73% for GPT-2). Note that GPT-2 trains with GPU exclaves, while VGG9 trains with CPU exclaves. Thus, the GPT2 training task occupies a smaller proportion of the overall time.
The time spent on executing the aggregation and model update exclaves increases linearly with the model size, because the performed averaging/addition operations have linear time complexity. Note that the DP exclave has a worst-case quadratic time complexity, because the SVT-DP algorithm iterates over all model parameters (potentially multiple times) to calculate the noise and exclude parameters beyond the noise threshold in the iteration. The total time spent on the DP, aggregation, and model update exclaves is less than 1% for VGG9 and up to 11% for GPT-2.
The execution time for the FL framework, including the time spent on orchestration and communication, also increases linearly with model size. Frequent communication and model (de)serialization makes the FL framework the second largest contributor to VerifiableFL’s execution time: 17% and 3% for GPT-2 and VGG9, respectively. This overhead comes from VerifiableFL’s orchestration of tasks that execute in exclaves: it is necessary for VerifiableFL to deploy exclaves in separate CVMs to provide the fine-grained runtime attestation of each exclave done by the hardware.
In addition, we evaluate the overhead of VerifiableFL’s data sanitization exclave, which a data provider deploys and runs prior to FL training on an Azure NCC40ads_H100_v5 CVM. The data sanitization exclave uses a classification model for harmful text, granite-guardian-hap-38m (granite), to detect toxic or abusive content and then removes the identified content to sanitize the dataset.
We measure the execution time of the data sanitization exclave when sanitizing each data provider’s Wikitext103 dataset. The result shows that the data sanitization exclave takes 257 seconds to sanitize each data owner’s dataset. This adds 2% overhead to VerifiableFL’s GPT-2 training. We conclude that enforce datasets properties using standard sanitisation techniques therefore is feasible in practice.
7.5. Overhead of runtime attestation
In this micro-benchmark, we measure the overhead due to runtime attestation. We use a single CVM with 4 vCPUs and 16 GB of memory to generate a SHA256 digest over 1 GB of input data and include the resulting hash digest as the runtime claim in an attestation request.
We observe that, on average, creating and signing an attestation report with vTPM takes 4 ms. This step has constant time complexity, because both creation and signing operations take a fixed-size hash digest as input. Generating hash digests is the most time-consuming operation during attestation: it takes 687 ms to calculate a SHA-256 hash over the 1 GB input data, and the time grows linearly with the input data size.
We also measure the attestation overhead with AMD’s PSP: the CVM requires less than 7 ms to fetch a signed attestation report from the PSP. Although we use vTPM attestation to emulate PSP attestation, the performance difference between them is minimal (under 3 ms per attestation) and negligible compared to VerifiableFL’s end-to-end training time, which takes hours.
(a) Throughput
(b) Latency
Figure 9. Read performance from secure storage
7.6. Overhead of secure storage
We also explore the performance impact of VerifiableFL’s secure storage for datasets. We mount a disk protected by dm-verity on an Azure DC32as_v5 CVM (P30 data disk with up to 200 MB/s and 5,000 IOPS). We then use the FIO benchmark (fio), which sequentially reads data from the ext-4 file system on the secure disk. We compare against insecure storage as a baseline, which uses a standard D32as_v5 VM on an unprotected Azure P30 disk.
As Fig. 9 shows, the sequential read throughput of the integrity-protected disk is 62% lower compared to regular storage. This is due to the cost of checking the block against the Merkle hash tree. Latency also increases accordingly: the secure storage module adds an additional $1.7\times$ overhead over the baseline latency (reported latencies include queuing time from the in-flight I/Os). The resulting read amplification impacts both throughput and latency, which adds an end-to-end performance cost to VerifiableFL.
Table 3. Overhead of auditing
#clients $\times$ Signature EDG total storage
#rounds verification (ms) creation (ms) (ms) (MB)
10 $\times$ 10 26 3 29 ¡1
10 $\times$ 100 196 13 209 2
10 $\times$ 1,000 1,886 131 2,017 17
100 $\times$ 10 218 20 238 2
100 $\times$ 100 1,802 146 1,948 16
100 $\times$ 1,000 17,404 1,837 19,241 162
7.7. Overhead of auditing
We evaluate the overhead that the auditing process incurs in real-world settings, when the number of clients scales from 10 to 100, and the number of rounds scales from 10 to 1,000. We generate synthetic EDRs and measure the auditor’s computation and storage overheads in a standard VM with 4 vCPUs and 16 GB memory.
Tab. 3 shows the results. When 100 clients train for 1000 rounds with VerifiableFL, it takes the auditor 19 seconds to verify all EDR signatures and construct the EDG, which is negligible in comparison to the FL training time. The incurred storage cost during auditing is 162 MB, which is also small. We conclude that the auditing process of VerifiableFL only introduces a small overhead, even with larger number of clients and training rounds.
8. Related Work
Mitigating attacks in FL. Various aggregation protocols are designed to mitigate privacy and integrity attacks in FL. Secure aggregation (bonawitz2017secure; sharma2019secure) addresses FL privacy concerns when aggregators are untrusted through cryptographic means. For example, Bonawitz et al. (bonawitz2017secure) implement secure aggregation using pair-wise Diffie-Hellman key agreement and t-out-of-n Shamir’s secret sharing, whilst Sharma et al. (sharma2019secure) use additive secret sharing. These protocols are complementary to VerifiableFL because they ensure the privacy of the FL training process, while VerifiableFL provides verifiable integrity guarantees that the correct computation was performed. Secure aggregation protocols could be used with VerifiableFL to provide both privacy and integrity guarantees.
VerifiableFL is compatible with a range of FL-specific techniques: robust aggregation (pillutla2022robust) make the FL training process more resilient against integrity attacks; Tolpegin et al. (tolpegin2020data) detect data poisoning attacks using outlier detection. To protect FL training against model poisoning attacks, Krum (blanchard2017machine) and Trimmed-mean (yin2018byzantine) use Byzantine-robust aggregation protocols for secure aggregation. While such protocols require at least a majority of trusted data providers, VerifiableFL’s verifiable claims hold with any number of malicious participants.
Trusted execution for FL. Mo et al. (mo_ppfl_2021) assume that both the FL server and clients are honest-but-curious. They use TEEs to prevent participants from observing the model data during training, thus preventing inference or data reconstruction attacks. Messaoud et al. (valerio21gradsec) address FL client-side privacy threats by protecting gradient updates inside TEEs. Since a subset of the model layers is prevented from gradient leakage, this mitigates against inference attacks. Both of the approaches assume a weaker threat model, without VerifiableFL’s detection of integrity attacks.
CrowdGuard (rieger_close_2022) uses TEEs to achieve integrity-preserving FL. At each training step, data providers use TEEs to evaluate the locally-trained models of every other data provider on their own data and vote regarding poisoned model. Unlike VerifiableFL, CrowdGuard assumes that TEEs offer confidentiality, which makes it vulnerable to side-channel attacks.
Ensuring integrity. Proof-carrying code (necula1997proof) and model checking (clarke1997model) can show that code complies with integrity policies by verifying it against a specification. Since VerifiableFL relies on execution integrity, it can be combined with such approaches.
Other approaches enforce control-flow integrity during execution: C-FLAT (abera2016cflat) and LO-FAT (lo-fat) attest intra-program control flow to ascertain the current execution state of a single node. In contrast, VerifiableFL attests task-level data transformations across distributed participants, allowing for verifiable claims about a final trained FL model. While there are similarities (e.g., use of hashes), the attestation objective is different: in particular, two FL training runs, despite identical control flow attestations, would still produce different models if the datasets are different.
9. Conclusions
We described VerifiableFL, a new system for executing FL jobs, while enabling an auditor to verify claims about the final trained FL model. VerifiableFL introduces exclaves as a new abstraction for the execution of FL tasks with integrity only, which attest all performed data transformations at runtime. Since executed FL tasks are attested through exclave data records (EDRs), it becomes impossible for participants to deviate from the correct FL training protocol, thus violating model claims, without detection. Exclaves produce EDRs, which can be combined to create an exclave dataflow graph (EDG) that spans the entire FL job. An auditor uses the EDG to verify end-to-end claims about the FL job. We showed that VerifiableFL can be integrated with a state-of-the-art FL framework to detect a wide range of deviations with low runtime overhead.